### v1

@@ 1 Introduction
One driving motivation behind the development of probabilistic appraoches to FEM is their potential application to (Bayesian) inverse problems. The typical procedure to infer a set of parameters of the PDE from a set of observations involves Monte Carlo sampling, with each sample requiring an FEM solve. To keep this approach computationally feasible, it might be necessary to resort to a coarse finite element discretization. However, this will reduce the accuracy of both the FEM solution and the resulting parameter estimates. The aim of taking a probabilistic approach to FEM is not necessarily to improve the accuracy of the posterior mean, but rather to incorporate the uncertainty due to discretization error in the posterior variance. We say that a probabilistic formulation of FEM is performing well in an inverse setting if 1) its posterior mean is at least as close to the ground truth as the posterior obtained with traditional FEM, and 2) its posterior variance accurately reflects the mismatch between the posterior mean and ground truth due to discretization error. The aim of this work is to investigate the performance of RM-FEM in the context of a Bayesian inverse problem, focussing specifically on the effects of interpolation error and solver bias.
@@ 3.1 Fixed observation nodes
3.2 2D and 3D problems
It is also worth noting that this small relative nodal error $\zeta$ is typical for FEM solutions to elliptic PDEs in 1D, but not in 2D or 3D. Additionally, RM-FEM has a different perturbation scheme in higher dimensions, sampling node locations from a uniform distribution over a disk or ball instead of an interval. These two factors could make the issue of posterior deterioration for the perturbed meshes not as severe for higher-dimensional PDEs, compared to 1D.
@@
To test this hypothesis, we change to a 2D domain $\Omega=(0,1)\times(0,\nicefrac{{1}}{{10}})$, with a Dirichlet boundary on the left and right $\Gamma_{d}=\{0,1\}\times(0,\nicefrac{{1}}{{10}})$ and a Neumann boundary on the top and bottom $\Gamma_{n}=(0,1)\times\{0,\nicefrac{{1}}{{10}}\}$, such that $\Gamma_{d}\cup\Gamma_{n}=\partial\Omega$ and $\Gamma_{d}\cap\Gamma_{n}=\emptyset$. We adapt the PDE such that $\kappa(x)$, $f(x)$ and $u(x)$ are all constant with respect to the vertical coordinate $y$, and the reference solution ${u}^{*}$ as a function of $x$ is exactly the same as the solution to Eq. 1:
$-\nabla\cdot\left(\kappa(x)\nabla u(x)\right)$ $=f(x)$ $\forall{\bm{x}}$ $\in\Omega$ (14)
$u$ $=0$ $\forall{\bm{x}}$ $\in\Gamma_{d}$
$\frac{\partial u}{\partial y}$ $=0$ $\forall{\bm{x}}$ $\in\Gamma_{n}$
@@
The coarsest mesh consists of 10 square elements with side length $h=\nicefrac{{1}}{{10}}$ and linear shape functions. Each finer mesh is obtained by splitting each element into 4 smaller square elements. For our RM-FEM implementation, we perturb each node by choosing from a uniform distribution on a disk with radius $r=\nicefrac{{1}}{{4}}\sqrt{2}h$, centered on the node location on the reference mesh [^1] Setting $r=\nicefrac{{1}}{{2}}h$ results in concave elements, which is why we use $r=\nicefrac{{1}}{{4}}\sqrt{2}h$ instead.. The boundary nodes are perturbed in the same manner, but projected back onto the boundary $\partial\Omega$ to ensure that the geometry of the domain is identical for each mesh. We define the observation locations for this 2D problem to be ${\bm{X}}=\{(\nicefrac{{i}}{{5}},\nicefrac{{1}}{{20}})\}_{i=1}^{4}$, and use the same observation values ${\bm{y}}$ as in the 1D case. When solving the inverse problem with FEM, we obtain exactly the same posterior distribution as in the 1D case, shown in Fig. 3(a).
@@
However, when using RM-FEM for this 2D case, we obtain the posterior shown in Fig. 5(b), which does differ from the 1D RM-FEM posterior shown in Fig. 3(b). Similarly to 1D RM-FEM with fixed observation nodes, we see in Table 1 that for this 2D problem, the width of the RM-FEM posterior has reduced significantly compared to 1D RM-FEM, and that the increase of confidence with mesh refinement is maintained. Furthermore, the error of the 2D RM-FEM posterior mean is smaller than that of 1D RM-FEM for all parameters $\xi_{i}$. Despite the difference between 1D and 2D RM-FEM not being as great as the difference between 1D RM-FEM with and without fixed observation nodes, we can conclude that both the deterioration of the posterior mean and the underconfidence are not as severe for higher-dimensional PDEs as for 1D problems.
@@ 4 Bias
Although we have seen in the previous section that the adverse effects of interpolation error on the performance of RM-FEM can be dampened to some extent, some points remain unexplained. For instance, throughout all examples shown in Figs. 3 and 5, $\xi_{1}$ can be seen to be consistently underestimated. As we will see, the reason for this lies not in some local effect like the one we explored in the previous section, but rather lies in a global bias in the FEM estimate of the solution.
@@
An additional observation to make about Fig. 6 is that the energy estimate provided by FEM is close to the optimal energy of the RM-FEM distribution. This is in line with common knowledge that smooth 1D differential equations like the one studied here are best discretized by a uniform mesh. More broadly, the aim of discretization strategies is to choose a mesh such that ${E}^{\text{h}}$ is a close to ${E}^{*}$ as possible for a given number of nodes or elements. A fundamental limitation of RM-FEM is that it pays little heed to the care that is typically put into mesh selection. If the original FEM mesh was chosen arbitrarily, then the RM-FEM mean energy $\mathbb{E}\left[{{\tilde{E}}}^{\text{h}}\right]$ would on average coincide with the FEM estimate ${E}^{\text{h}}$. However, since the FEM mesh is chosen close to optimally, we see a deterioration of the RM-FEM mean energy $\mathbb{E}\left[{{\tilde{E}}}^{\text{h}}\right]$ when comparing against the FEM estimate ${E}^{\text{h}}$.
@@ 5 Conclusion
In this work, we investigated the effects of interpolation error and bias on the performance of RM-FEM in an inverse setting. In 1D, interpolation error can deteriorate the method’s performance, yielding a posterior mean that is further away from the ground truth and an overly large variance, when compared to the posterior mean that is obtained using traditional FEM. This effect can be dampened by excluding the nodes at the observation locations from the perturbation scheme, and is generally less pronounced in 2D settings. A more pernicious drawback of RM-FEM is its quantification of discretization error for quantities of interest for which FEM is known to be biased, in which case the RM-FEM distribution is biased as well, possibly even more so. Although RM-FEM makes a useful tool for error estimation and adaptive mesh refinement, we recommend caution when employing it in an inverse setting, particularly in cases where FEM produces biased estimates of the observed quantity of interest.

### v2

@@ 1 Introduction
One driving motivation behind the development of probabilistic approaches to FEM is their potential application to (Bayesian) inverse problems. The typical procedure to infer a set of parameters of the PDE from a set of observations involves Monte Carlo sampling, with each sample requiring an FEM solve. To keep this approach computationally feasible, it might be necessary to resort to a coarse finite element discretization. However, this will reduce the accuracy of both the FEM solution and the resulting parameter estimates. The aim of taking a probabilistic approach to FEM is not necessarily to improve the accuracy of the posterior mean, but rather to incorporate the uncertainty due to discretization error in the posterior variance. We say that a probabilistic formulation of FEM is performing well in an inverse setting if 1) its posterior mean is at least as close to the ground truth as the posterior obtained with traditional FEM, and 2) its posterior variance accurately reflects the mismatch between the posterior mean and ground truth due to discretization error. The aim of this work is to investigate the performance of RM-FEM in the context of a Bayesian inverse problem, focussing specifically on the effects of interpolation error and solver bias.
@@ 3.1 Fixed observation nodes
3.2 Generalizability
We note that the conclusion of the previous section is quite specific to the 1D case with nodal observations. If observations are made over a small region rather than a single point, the effect of interpolation error would be dampened. Similarly, this relatively large interpolation error $\eta$ is typical for FEM solutions to elliptic PDEs in 1D, but not in 2D or 3D. For more realistic problems, most of the discretization error is due to nodal error $\zeta$ rather than interpolation error $\eta$. Finally, higher-dimensional problems require a different perturbation scheme to guarantee a valid finite element discretization. For example, in the case of quadrilateral elements, the maximal perturbation that still guarantees mesh validity is a disk centered on the nodes with a radius $r=\nicefrac{{1}}{{4}}\sqrt{2}h$. Considering only the $x$-component, we would obtain a semicircular distribution on the support $(-\nicefrac{{1}}{{4}}\sqrt{2}h,\nicefrac{{1}}{{4}}\sqrt{2}h)$ rather than a uniform distribution on the support $(-\nicefrac{{1}}{{2}}h,\nicefrac{{1}}{{2}}h)$.
@@ 3.2 Generalizability
Although all three of these factors are expected to reduce the relative effect of interpolation error in a more realistic setting, we will single out the final point (i.e., a different perturbation scheme) to put it to the test. We change to a 2D domain $\Omega=(0,1)\times(0,\nicefrac{{1}}{{10}})$, with a Dirichlet boundary on the left and right $\Gamma_{d}=\{0,1\}\times(0,\nicefrac{{1}}{{10}})$ and a Neumann boundary on the top and bottom $\Gamma_{n}=(0,1)\times\{0,\nicefrac{{1}}{{10}}\}$, such that $\Gamma_{d}\cup\Gamma_{n}=\partial\Omega$ and $\Gamma_{d}\cap\Gamma_{n}=\emptyset$. We adapt the PDE such that $\kappa({\bm{x}})$, $f({\bm{x}})$ and $u({\bm{x}})$ are all constant with respect to the vertical coordinate $y$, and the reference solution ${u}^{*}$ as a function of $x$ is exactly the same as the solution to Eq. 1:
$-\nabla\cdot\left(\kappa({\bm{x}})\nabla u({\bm{x}})\right)$ $=f({\bm{x}})$ $\forall{\bm{x}}$ $\in\Omega$ (14)
$u({\bm{x}})$ $=0$ $\forall{\bm{x}}$ $\in\Gamma_{d}$
$\frac{\partial u}{\partial y}({\bm{x}})$ $=0$ $\forall{\bm{x}}$ $\in\Gamma_{n}$
@@
The coarsest mesh consists of 10 square elements with side length $h=\nicefrac{{1}}{{10}}$ and linear shape functions. Each finer mesh is obtained by splitting each element into 4 smaller square elements. For our RM-FEM implementation, we perturb each node by choosing from a uniform distribution on a disk with radius $r=\nicefrac{{1}}{{4}}\sqrt{2}h$, centered on the node location on the reference mesh. The boundary nodes are perturbed in the same manner, but projected back onto the boundary $\partial\Omega$ to ensure that the geometry of the domain is identical for each mesh. We define the observation locations for this 2D problem to be ${\bm{X}}=\{(\nicefrac{{i}}{{5}},\nicefrac{{1}}{{20}})\}_{i=1}^{4}$, and use the same observation values ${\bm{y}}$ as in the 1D case. When solving the inverse problem with FEM, we obtain exactly the same posterior distribution as in the 1D case, shown in Fig. 3(a).
@@
However, when using RM-FEM for this 2D case, we obtain the posterior shown in Fig. 5(b), which does differ from the 1D RM-FEM posterior shown in Fig. 3(b). Similarly to 1D RM-FEM with fixed observation nodes, we see in Table 1 that for this 2D problem, the width of the RM-FEM posterior has reduced significantly compared to 1D RM-FEM, and that the increase of confidence with mesh refinement is maintained. Furthermore, the error of the 2D RM-FEM posterior mean is smaller than that of 1D RM-FEM for all parameters $\xi_{i}$. Despite the difference between 1D and 2D RM-FEM not being as great as the difference between 1D RM-FEM with and without fixed observation nodes, we can conclude that the 2D perturbation scheme reduces the severity of the deterioration of the posterior mean. We expect that the deterioration of the posterior mean would vanish altogether for problems with regional observations (as opposed to point observations) and problems where the relative nodal error $\zeta$ dwarfs the relative interpolation error $\eta$.
@@ 4 Bias
Although we have seen in the previous section that the adverse effects of interpolation error on the performance of RM-FEM can be dampened to some extent, some points remain unexplained. In particular, it is still unclear why RM-FEM does not necessarily improve the accuracy of the parameter estimate provided by the posterior mean compared to FEM. For instance, throughout all examples shown in Figs. 3 and 5, $\xi_{1}$ can be seen to be consistently underestimated. As we will see, the reason for this lies not in some local effect like the one we explored in the previous section, but rather lies in a global bias in the FEM estimate of the solution.
@@
An additional observation to make about Fig. 6 is that the energy estimate provided by FEM is close to the optimal energy of the RM-FEM distribution. This is in line with common knowledge that smooth 1D differential equations like the one studied here are best discretized by a uniform mesh. More broadly, the aim of discretization strategies is to choose a mesh such that ${E}^{\text{h}}$ is as close to ${E}^{*}$ as possible for a given number of nodes or elements. A fundamental limitation of RM-FEM is that it pays little heed to the care that is typically put into mesh selection. If the original FEM mesh was chosen arbitrarily, then the RM-FEM mean energy $\mathbb{E}\left[{{\tilde{E}}}^{\text{h}}\right]$ would on average coincide with the FEM estimate ${E}^{\text{h}}$. However, since the FEM mesh is chosen close to optimally, we see a deterioration of the RM-FEM mean energy $\mathbb{E}\left[{{\tilde{E}}}^{\text{h}}\right]$ when comparing against the FEM estimate ${E}^{\text{h}}$.
@@ 5 Conclusion
In this work, we investigated the effects of interpolation error and bias on the performance of RM-FEM in an inverse setting. In 1D, interpolation error can deteriorate the method’s performance, yielding a posterior mean that is further away from the ground truth and an overly large variance, when compared to the posterior mean that is obtained using traditional FEM. This effect can be dampened by excluding the nodes at the observation locations from the perturbation scheme, and is generally less pronounced in more practical settings. A more pernicious drawback of RM-FEM is its quantification of discretization error for quantities of interest for which FEM is known to be biased, in which case the RM-FEM distribution is biased as well, possibly even more so. Although RM-FEM makes a useful tool for error estimation and adaptive mesh refinement, we recommend caution when employing it in an inverse setting, particularly in cases where FEM produces biased estimates of the observed quantity of interest.